### v1

# Mechanically concealed holes

Abstract
When a hole is introduced into an elastic material, it will usually act to reduce the overall mechanical stiffness. A general ambition is to investigate whether a stiff shell around the hole can act to maintain the overall mechanical properties. We consider the basic example situation of an isotropic, homogeneous, linearly elastic material loaded uniformly under plane strain for low concentrations of holes. As we demonstrate, the thickness of the shell can be adjusted in a way to maintain the overall stiffness of the system. We derive a corresponding mathematical expression for the thickness of the shell that conceals the hole. Thus, one can work with given materials to mask the presence of the holes. One does not necessarily need to adjust the material parameters and thus materials themselves. Our predictions from linear elasticity continuum theory are extended to atomistic level using molecular dynamics simulations of a model Lennard-Jones solid. Small deviations from linear elasticity theory can be minimized by tuning the hole-to-system size ratio in the molecular dynamics simulations. This extension attests the robustness of our continuum predictions even at atomistic scales. The basic concept is important in the context of light-weight construction.
Introduction. Saving resources and fuel is a major concern of recent production lines and construction design. Besides aspects of sustainability, pure economic reasons favor corresponding achievements. Saving materials and energy reduces overall costs [1]. Therefore, light-weight construction remains key to future technological developments. Nature provides corresponding examples, maybe bones being the most obvious ones [2, 3]. Thanks to their stiff structure they provide overall stability for the whole organism, yet the many cavities of various types of bones reduce their overall weight. This structure saves energy of motion and increases agility and mobility.
Analogously, for many components of machines, vehicles, aircrafts, or other devices, reducing weight provides significant benefit [4, 5]. However, the overall design has often been developed for years or decades and been adjusted to near perfection. In such cases, changing the dimension or shape of individual components to reduce their weight, or others of their mechanical properties like stiffness, provides additional challenges.
Therefore, we focus on the idea of introducing holes into materials to save weight [6, 7]. Advanced strategies of designing materials with cylindrical or spherical holes had been already found beneficial for this purpose using approaches of topological optimization [8]. In our case, holes are considered to be mechanically masked in a way so that their presence is not noted on the overall, macroscopic scale of the material. Still, the overall mechanical stiffness shall be maintained. Together, this concept results in a component of identical mechanical properties, yet of reduced weight. Key is to introduce holes (cavities) that are surrounded by stiffer shells so that the overall, combined mechanical stiffness is the same as in the absence of the holes.
The model geometry of introducing hollow cylindrical shells in three-dimensional solids (or hollow circular rings in two-dimensional solids) has also been applied to study diverse other problems, ranging from elastostatics [9, 10] to cavitation in soft solids [11, 12]. Amongst these, the idea of mechanically masking holes inside a solid, in fact, has been studied before in terms of “mechanical cloaking”. It provides the solid with a property of “mechanical unfeelability” of the holes within, in terms of the overall macroscopic response. Additionally, mechanically masking holes and corresponding optimization has been realized to conceal mechanical [13], thermo-mechanical [14], as well as dispersive properties [15] of solids, especially in the context of metamaterials design [16, 17, 18]. Linear elastodynamic cloaking of cylindrical holes was discussed for infinitesimal in-plane deformations [19].
Recently, mechanical concealment in elastostatic situations received additional attention. Notably, when viewed from the perspective of optimized materials design, elastostatic cloaking has been outlined following two strategies. Either the displacement fields [20] or the elastic moduli [21] and thus the type of employed materials were used as design parameters. Interestingly, a recent analytical and numerical work [22] showed that elastostatic mechanical cloaking of a circular inclusion in two-dimensional geometries can be attained by coating the inclusion using several concentrically arranged circular rings and tuning their shear moduli. This strategy eliminates the complexity of additionally tuning the Poisson ratio.
Yet, in reality, the materials to be used may already be determined by other factors, such as constraints of production processes or cost. Prescribing the materials to be used also fixes the elastic moduli, so that these parameters can hardly be adjusted. In that case, a different recipe to implement mechanical shielding and concealing of holes in a given solid is necessary.
To this end, we present an alternative strategy to illustrate the concept of mechanically concealed holes using a basic example situation. We address uniform loading under plane-strain conditions of a material that is homogeneous, isotropic, linearly elastic, and, in principle, infinitely extended. The cylindrical holes that we introduce are of sufficiently low concentration so that we may neglect their mutual mechanical interaction. In reality, we may, for instance, think of cylindrical holes drilled into plates or blocks of material. Using analytical calculations as well as molecular dynamics simulations, we demonstrate that, indeed, under such conditions, we can introduce stiff cylindrical shells around the cylindrical holes so that the presence of the holes is effectively masked and mechanically concealed. Our sole tuning parameter is the thickness of the shell surrounding the hole.
Theoretical background. We start our theoretical consideration from the basic theory of linear elasticity [23]. Stress $\bm{\sigma}^{\mathrm{(3d)}}$ and strain $\bm{\varepsilon}^{\mathrm{(3d)}}$ are related to each other via the shear modulus $\mu$ and the Lamé parameter $\lambda$, which is associated with compressibility,
$\bm{\sigma}^{\mathrm{(3d)}}=2\mu\bm{\varepsilon}^{\mathrm{(3d)}}+\lambda\mathbf{I}\varepsilon_{kk}^{\mathrm{(3d)}}.$ (1)
Einstein summation convention is applied and $\mathbf{I}$ denotes the unit matrix. The second Lamé parameter can be expressed by $\mu$ and the Poisson ratio $\nu$ via $\lambda=2\mu\nu/(1-2\nu)$.
In our plane-strain geometry, we denote in Cartesian coordinates the plane as spanned by coordinates $x$ and $y$, while the direction normal to the plane is referred to by the coordinate $z$. Consequently, plane-strain conditions imply $\varepsilon_{xz}^{\mathrm{(3d)}}=\varepsilon_{zx}^{\mathrm{(3d)}}=\varepsilon_{yz}^{\mathrm{(3d)}}=\varepsilon_{zy}^{\mathrm{(3d)}}=\varepsilon_{zz}^{\mathrm{(3d)}}=0$. Thus, solving Eq. (1) for $\sigma_{zz}^{\mathrm{(3d)}}$, we can express the whole remaining physics in the two-dimensional plane in terms of the two-dimensional stress $\bm{\sigma}$ and strain $\bm{\varepsilon}$ as
$\bm{\varepsilon}=\frac{1}{2\mu}\left(\bm{\sigma}-\nu\mathbf{I}\sigma_{kk}\right)$ (2)
and
$\bm{\nabla}\cdot\bm{\sigma}=\bm{0}.$ (3)
The latter condition is directly satisfied by deriving $\bm{\sigma}$ from the stress function $F$ as
$\bm{\sigma}=\mathbf{I}\bm{\nabla}^{2}F-\bm{\nabla}\bm{\nabla}F.$ (4)
If the associated strain $\bm{\varepsilon}$ derives from a displacement field $\mathbf{u}$, the strain must satisfy certain compatibility conditions. They ensure that our strain can be expressed as
$\bm{\varepsilon}=\frac{1}{2}\left(\bm{\nabla}\mathbf{u}+\left(\bm{\nabla}\mathbf{u}\right)^{T}\right),$ (5)
where T marks the transpose. Specifically, in our two-dimensional plane-strain geometry, these compatibility relations reduce to
$\nabla_{1}\nabla_{1}\varepsilon_{22}+\nabla_{2}\nabla_{2}\varepsilon_{11}-2\nabla_{1}\nabla_{2}\varepsilon_{12}=0.$ (6)
The subscripts 1 and 2 mark two orthogonal coordinates in the two-dimensional plane. From Eq. (2) we can calculate the corresponding strain associated with the two-dimensional stress given by Eq. (4). Inserting it into Eq. (6) implies that compatibility with Eq. (5) is ensured, if
$\bm{\nabla}^{2}\bm{\nabla}^{2}F=0.$ (7)
Since Eq. (3) is automatically satisfied in this case, the condition in Eq. (7) is sufficient and necessary for this solution to exist.
Derivation of the mechanical solution for a shelled hole. We address the situation in polar coordinates. Our coordinate system is centered in the hole, see Fig. 1. The hole has a radius $a$, while the surrounding shell of inner radius $a$ is of outer radius $b>a$. While the shell is of shear modulus $\mu^{\mathrm{i}}$ and Poisson ratio $\nu^{\mathrm{i}}$, the corresponding parameters of the surrounding elastic material are $\mu^{\mathrm{o}}$ and $\nu^{\mathrm{o}}$. Here and in the following, superscript “i” indicates the “inner” elastic material (shell), while “o” denotes the “outer” elastic material.
Figure 1: Illustration of the geometry. In the two-dimensional plane that we use to describe the block of material under plane-strain conditions, the cylindrical hole appears as a circular exclusion of radius $a$. Our system of polar coordinates $(r,\varphi)$ is centered in the hole. The hole is surrounded by a cylindrical shell of outer radius $b$, mechanical shear modulus $\mu^{\mathrm{i}}$, and Poisson ratio $\nu^{\mathrm{i}}$. Moreover, the actual, outer elastic material is of shear modulus $\mu^{\mathrm{o}}$ and Poisson ratio $\nu^{\mathrm{o}}$.
Since we consider uniform loading, the stress tensor at infinite distance from the hole can be denoted as $\bm{\sigma}(r\rightarrow\infty)=-P\mathbf{I}$, where $P>0$. In this circularly symmetric situation, for linearly elastic systems, there is no angular dependence of the results on the polar angle $\varphi$. Thus, from all possible terms contributing to the stress function $F$ that satisfy Eq. (7) [24], we retain only those that do not imply any dependence on $\varphi$ in the physical solution. We formulate the stress function separately for the inner and outer regions,
$F^{\{\mathrm{i},\mathrm{o}\}}=A^{\{\mathrm{i},\mathrm{o}\}}r^{2}+B^{\{\mathrm{i},\mathrm{o}\}}\varphi+C^{\{\mathrm{i},\mathrm{o}\}}\ln(r).$ (8)
In polar coordinates,
$\sigma_{rr}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{1}{r}\frac{\partial F^{\{\mathrm{i},\mathrm{o}\}}}{\partial r}+\frac{1}{r^{2}}\frac{\partial^{2}F^{\{\mathrm{i},\mathrm{o}\}}}{\partial\varphi^{2}},$ (9)
$\sigma_{\varphi\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{\partial^{2}F^{\{\mathrm{i},\mathrm{o}\}}}{\partial r^{2}},$ (10)
$\sigma_{r\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\sigma_{\varphi r}^{\{\mathrm{i},\mathrm{o}\}}={}-\frac{1}{r}\frac{\partial^{2}F^{\{\mathrm{i},\mathrm{o}\}}}{\partial r\partial\varphi}+\frac{1}{r^{2}}\frac{\partial F^{\{\mathrm{i},\mathrm{o}\}}}{\partial\varphi}.$ (11)
Thus, Eq. (4) leads us to
$\sigma_{rr}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $2A^{\{\mathrm{i},\mathrm{o}\}}+C^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}},$ (12)
$\sigma_{\varphi\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $2A^{\{\mathrm{i},\mathrm{o}\}}-C^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}},$ (13)
$\sigma_{r\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\sigma_{\varphi r}^{\{\mathrm{i},\mathrm{o}\}}=B^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}}.$ (14)
From Eq. (2), we find the expressions for the components of the strain tensor
$\epsilon_{rr}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{1}{2\mu^{\{\mathrm{i},\mathrm{o}\}}}\Bigg(2\left(1-2\nu^{\{\mathrm{i},\mathrm{o}\}}\right)A^{\{\mathrm{i},\mathrm{o}\}}+C^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}}\Bigg),\qquad$ (15)
$\epsilon_{\varphi\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{1}{2\mu^{\{\mathrm{i},\mathrm{o}\}}}\Bigg(2\left(1-2\nu^{\{\mathrm{i},\mathrm{o}\}}\right)A^{\{\mathrm{i},\mathrm{o}\}}-C^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}}\Bigg),$ (16)
$\epsilon_{r\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\epsilon_{\varphi r}^{\{\mathrm{i},\mathrm{o}\}}=\frac{1}{2\mu^{\{\mathrm{i},\mathrm{o}\}}}B^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r^{2}}.$ (17)
In polar coordinates, the relations between the strain and displacement field
$\epsilon_{rr}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{\partial u_{r}^{\{\mathrm{i},\mathrm{o}\}}}{\partial r},$
$\epsilon_{\varphi\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{1}{r}\left(\frac{\partial u_{\varphi}^{\{\mathrm{i},\mathrm{o}\}}}{\partial\varphi}+u_{r}^{\{\mathrm{i},\mathrm{o}\}}\right),$
$\epsilon_{r\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\epsilon_{\varphi r}^{\{\mathrm{i},\mathrm{o}\}}=\frac{1}{2}\left(\frac{1}{r}\frac{\partial u_{r}^{\{\mathrm{i},\mathrm{o}\}}}{\partial\varphi}+\frac{\partial u_{\varphi}^{\{\mathrm{i},\mathrm{o}\}}}{\partial r}-\frac{u_{\varphi}^{\{\mathrm{i},\mathrm{o}\}}}{r}\right)\qquad$ (18)
apply. From here, we obtain the displacement fields
$u_{r}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ $\frac{1}{2\mu^{\{\mathrm{i},\mathrm{o}\}}}\Bigg(2\left(1-2\nu^{\{\mathrm{i},\mathrm{o}\}}\right)A^{\{\mathrm{i},\mathrm{o}\}}r-C^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r}\Bigg),\qquad$ (19)
$u_{\varphi}^{\{\mathrm{i},\mathrm{o}\}}$ $=$ ${}-\frac{1}{2\mu^{\{\mathrm{i},\mathrm{o}\}}}B^{\{\mathrm{i},\mathrm{o}\}}\frac{1}{r}.\qquad$ (20)
Our next task is to obtain the values of the coefficients $A^{\{\mathrm{i},\mathrm{o}\}}$, $B^{\{\mathrm{i},\mathrm{o}\}}$, and $C^{\{\mathrm{i},\mathrm{o}\}}$ from the boundary conditions. First, the surface of the hole must be free of traction forces,
$\sigma_{rr}^{\mathrm{i}}(r=a)$ $=$ $0,$ (21)
$\sigma_{r\varphi}^{\mathrm{i}}(r=a)$ $=$ $0.$ (22)
Next, at infinite distance from the hole, the stress must be of the imposed form
$\bm{\sigma}^{\mathrm{o}}(r\rightarrow\infty)={}-P\mathbf{I}.$ (23)
At the interface between the shell and the surrounding elastic material, both radial stress components must match each other,
$\sigma_{rr}^{\mathrm{i}}(r=b)$ $=$ $\sigma_{rr}^{\mathrm{o}}(r=b),$ (24)
$\sigma_{r\varphi}^{\mathrm{i}}(r=b)$ $=$ $\sigma_{r\varphi}^{\mathrm{o}}(r=b),$ (25)
as must the components of the displacement field,
$u_{r}^{\mathrm{i}}(r=b)$ $=$ $u_{r}^{\mathrm{o}}(r=b),$ (26)
$u_{\varphi}^{\mathrm{i}}(r=b)$ $=$ $u_{\varphi}^{\mathrm{o}}(r=b).$ (27)
From all these conditions, we find the magnitudes of the coefficients
$A^{\mathrm{i}}$ $=$ ${}-\frac{\left(1-\nu^{\mathrm{o}}\right)Pb^{2}}{\frac{\mu^{\mathrm{o}}}{\mu^{\mathrm{i}}}\left[\left(1-2\nu^{\mathrm{i}}\right)b^{2}+a^{2}\right]+b^{2}-a^{2}},$ (28)
$A^{\mathrm{o}}$ $=$ ${}-\frac{1}{2}P,$ (29)
$B^{\mathrm{i}}$ $=$ $B^{\mathrm{o}}={}0,$ (30)
$C^{\mathrm{i}}$ $=$ $\frac{2\left(1-\nu^{\mathrm{o}}\right)Pa^{2}b^{2}}{\frac{\mu^{\mathrm{o}}}{\mu^{\mathrm{i}}}\left[\left(1-2\nu^{\mathrm{i}}\right)b^{2}+a^{2}\right]+b^{2}-a^{2}},$ (31)
$C^{\mathrm{o}}$ $=$ ${}\left(1-\frac{2\left(1-\nu^{\mathrm{o}}\right)\left(b^{2}-a^{2}\right)}{\frac{\mu^{\mathrm{o}}}{\mu^{\mathrm{i}}}\left[\left(1-2\nu^{\mathrm{i}}\right)b^{2}+a^{2}\right]+b^{2}-a^{2}}\right)Pb^{2}.\qquad$ (32)
In this way, we have determined the expressions for the displacements, strains, and stresses in the entire domain. For identical materials of the shell around the hole and the surrounding elastic body, that is, for $\mu^{\mathrm{i}}=\mu^{\mathrm{o}}$ and $\nu^{\mathrm{i}}=\nu^{\mathrm{o}}$, we recover the solution for a uniform elastic body containing a hole, namely $A^{\mathrm{i}}=A^{\mathrm{o}}=-P/2$, $B^{\mathrm{i}}=B^{\mathrm{o}}=0$, and $C^{\mathrm{i}}=C^{\mathrm{o}}=Pa^{2}$.
Figure 2: Ratio $b/a$ between the outer radius of the shell $b$ around the hole and the radius $a$ of the hole as a function of the given ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ between the mechanical moduli of the shell and the surrounding elastic material. The value of $b$ is chosen in a way to mechanically conceal and mask the presence of the hole in the surrounding elastic substance under the imposed deformation. Curves are shown for different combinations of the Poisson ratios $1<\nu^{\mathrm{\{i,o\}}}<1/2$ of the shell and the surrounding elastic material, (a) both nonauxetic ($\nu^{\mathrm{\{i,o\}}}>0$) , (b) both auxetic ($\nu^{\mathrm{\{i,o\}}}<0$), and (c) combined auxetic and nonauxetic ($\nu^{\mathrm{i}}\nu^{\mathrm{o}}<0$).
Thickness of the shell for mechanical concealment. We now turn to the central point. The shell around the hole shall mechanically conceal and mask the hole in a way that its presence is not noted from outside. This concealment shall be realized for given materials, that is, we may not modify the material parameters $\mu^{\{\mathrm{i,o}\}}$ and $\nu^{\{\mathrm{i,o}\}}$. For that purpose, we must choose the outer radius $b$ of the shell in a way that the mechanical solution outside the shelled hole appears in the same way as if the hole were not present at all.
Specifically, this means that the stress outside the shell is given by the imposed uniform stress, $\bm{\sigma}^{\mathrm{o}}(r>b)=-P\mathbf{I}$. From Eqs. (12)–(14), together with Eq. (30), this implies
$C^{\mathrm{o}}=0.$ (33)
This condition further guarantees that also the strain $\bm{\varepsilon}^{\mathrm{o}}(r>b)$ and the displacement field $\mathbf{u}^{\mathrm{o}}(r>b)$ adapt their values in a uniform elastic material as if the hole and its shell were absent.
Indeed, Eqs. (32) and (33) can be solved for the outer radius $b$ of the shell,
$b=\sqrt{\frac{\frac{\mu^{\mathrm{i}}}{\mu^{\mathrm{o}}}\left(1-2\nu^{\mathrm{o}}\right)+1}{\frac{\mu^{\mathrm{i}}}{\mu^{\mathrm{o}}}\left(1-2\nu^{\mathrm{o}}\right)-\left(1-2\nu^{\mathrm{i}}\right)}}\,a.$ (34)
The Poisson ratios are confined to values $1<\nu^{\{\mathrm{i},\mathrm{o}\}}<1/2$. Thus, for shells sufficiently stiffer than the outer elastic material, $\mu^{\mathrm{i}}\gg\mu^{\mathrm{o}}$, the expression for $b$ always exists. Complete mechanical shielding of the hole in the surrounding elastic material by a stiff shell is therefore always possible under the given deformation. We illustrate $b$ as a function of the ratio of elastic shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for different combinations of Poisson ratios $\nu^{\mathrm{i}}$ and $\nu^{\mathrm{o}}$ in Fig. 2. For identical elastic materials ($\mu^{\mathrm{i}}=\mu^{\mathrm{o}}$), whether both of them are nonauxetic as in Fig. 2(a) or auxetic as in Fig. 2(b), the necessary outer radius of the shell and thus the necessary thickness of the shell bab-a diverge. In contrast to that, for very stiff shell materials $\mu^{\mathrm{i}}\gg\mu^{\mathrm{o}}$, the thickness bab-a of the shell tends towards zero ($b/a\rightarrow 1$ in Fig. 2).
If the Poisson ratio of the shell is smaller compared to the surrounding material, $\nu^{\mathrm{i}}<\nu^{\mathrm{o}}$, the necessary outer radius and thus thickness of the shell diverges even with shells stiffer than the outside elastic material, $\mu^{\mathrm{i}}>\mu^{\mathrm{o}}$, if the shells are not overly stiff, see the red curves in Fig. 2. In such cases, for given materials, mechanical shielding will be challenging, and a stiffer choice of shell material may still be advisable. For example, the combination $\nu^{\mathrm{i}}=0.3$ and $\nu^{\mathrm{o}}=0.49$, see Fig. 2(a), depicts a compressible shell such as steel [25, 26] within a nearly incompressible background like a rubbery elastomer [25, 26]. However, given that steels are generally several orders of magnitude stiffer than elastomers, this does not represent an actual constraint for mechanical concealment. It would still be possible already using very thin shells.
Interestingly, mechanical concealment would also be possible in the reverse situation of $\nu^{\mathrm{i}}>\nu^{\mathrm{o}}$ for a hard incompressible elastomeric shell surrounding a hole in a soft compressible solid, as long as the shell is not substantially softer than the surrounding solid, $\mu^{\mathrm{i}}\lessapprox\mu^{\mathrm{o}}$. This statement applies irrespectively of the absolute values of $\nu^{\mathrm{\{i,o\}}}$, see the blue dash-dotted lines in Fig. 2. Yet, mechanical concealment of the hole is only achieved at elevated thicknesses of the shell ($5<b/a<15$).
Thus, where possible, selecting most beneficial materials provides advantages. Leveraging the existence of various unconventional materials, one may even opt for materials of near-zero or negative Poisson ratios [27, 28, 29, 30, 31, 32, 33] to achieve advanced mechanical shielding. However, once a choice is made or the materials are determined by other factors, then still for many practical material combinations the strategy of simply adjusting the thickness of the shell provides an effective, realizable way for mechanical concealment.
Molecular dynamics simulations. To proceed a step beyond continuum considerations, we seek to investigate whether our analytical solution for mechanical concealment of a hole carries over to atomistic level. For this purpose, we analyze a typical microscopic model solid using molecular dynamics (MD) simulations. We set combinations of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ and $b/a$ and test whether mechanical concealment can be achieved.
In fact, atomistic investigations of mechanically concealed holes inside a solid are rare and challenging. Specifically, stabilizing solids with voids or cavities on the atomic scale often needs many-body terms in the interatomic interaction potentials. They come at a cost of losing simple, transparent models. Often, the descriptions invoke anisotropy due to angular dependence of the potentials. Our scope is to circumvent this issue and work with a description as simple as possible. Therefore, we use a two-dimensional, truncated and shifted Lennard-Jones (LJ) potential and adjust its length and energy scale, that is, the depth of the energy well, see the supplemental material for details [34], even if resulting energy wells are significantly deeper than for standard rare-gas solids. They imply very high mechanical moduli of the order of TPa [34]. As an advantage, we can work with a very simple and accessible potential and thus mimic a “toy model” for a strongly cohesive solid. As a further benefit, the LJ potential in two dimensions leads to stable hexagonal solids. The linearly elastic properties of solids of hexagonal symmetry are the same as for isotropic ones from a continuum perspective [23]. Therefore, our atomistic approach is in line with the continuum considerations above for isotropic elastic materials.
Figure 3: MD simulation results for the variation of pressure $P^{*}$ with areal strain $\epsilon_{A}=\Delta A/A_{0}$ during isotropic compression. Here, $A_{0}$ is the area of the equilibrated system at a given initial pressure $P^{*}(\epsilon_{A}=0)$, and $\Delta A$ is the change in area from there. We define the bulk moduli $K^{*}$ as the slopes of the resulting curves. Both pressure $P^{*}=P\sigma^{2}/\epsilon$ and bulk modulus $K^{*}=K\sigma^{2}/\epsilon$ are rescaled by the LJ parameters. We consider the two ratios of elastic moduli between the shell and the surrounding solid (a) $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ and (b) $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$. In both cases, $P^{*}(\epsilon_{A})$ is plotted for the pristine planar hexagonal solid without any hole (red curves), the solid with an unshelled hole ($b/a=1$, black curves), and for the solid with a shelled hole of a thickness $b/a>1$ (blue curves) that best ensure mechanical concealment (same slope $K^{*}\approx 54.98$ for red and blue curves). The corresponding geometries were identified as (c) $b/a=2.018$ and (d) $b/a=1.173$, respectively, in contrast to (e) the unshelled hole that yields a lower $K^{*}\approx 50.2$. The side panels show simulation snapshots with gray and blue atoms denoting whether they belong to the background solid or shell, respectively. We considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34 in all cases.
For a detailed overview on the parameter values and protocol of our MD simulations, we refer to the supplemental material [34], which includes the additional Refs. 35, 36, 37, 38, 39, 40, 41. We here just summarize that strongly cohesive, two-dimensional, planar solids consisting of 46200 atoms that interact via a truncated and force-shifted Lennard-Jones (LJ) potential were simulated under periodic boundary conditions. Very low temperatures ($T$ = 0.6 K) were imposed throughout to allow for comparison with the continuum analogs described above. The characteristic size of the atoms was set to $\sigma=1$ Å and the radius of the holes carved out from the solid to $a=20$ Å. Pristine hexagonal solids, hexagonal solids with a hole but no shell ($b/a=1$), and hexagonal solids containing shelled holes ($b/a>1$) were prepared. We here use the bulk modulus to compare their mechanical behavior and test the effective mechanical shielding of the hole from isotropic compression. To introduce stiff shells, we set the minimum of the LJ potential deeper for shell atoms than for atoms belonging to the surrounding elastic solid. At long wavelengths the LJ parameters for a two-dimensional hexagonal lattice can be related to the resulting effective shear modulus [42, 43]. We rely on this relation to define and obtain the ratio between our elastic moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ [34]. Our MD simulations were performed using LAMMPS [44, 45].
We proceed as follows to achieve mechanical concealment of the hole. Initially, we choose a given stiffness of the shell surrounding the hole by setting a specific value of the ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. This step corresponds to accepting the situation of given materials that we need to work with. Then we tune the thickness of the shell by suitably varying $b/a$. The desired $b/a$ is obtained by matching the bulk modulus $K$ of the hexagonal solid with shelled hole to that of the pristine planar hexagonal solid without hole. Figure 3 represents two examples of such numerical experiments, both for $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ = 0.34, yet for (a) $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ and (b) $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$. In both Figs. 3(a) and (b), the pristine hexagonal solid without any hole yields an elastic bulk modulus $K=54.98\epsilon/\sigma^{2}$ (red curves). When we cut out the hole, but do not put any shell surrounding it ($b/a=1$), see Fig. 3(e), the bulk modulus is reduced to $K=50.2\epsilon/\sigma^{2}$ (black curves). However, we can restore the initial bulk modulus of $K=54.98\epsilon/\sigma^{2}$ by placing a stiff shell of matched thickness enclosing the hole (blue curves). The corresponding thicknesses to conceal the presence of the hole are given by $b/a\approx 2.018$ for the softer shell material $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$, see Fig. 3(c), and $b/a\approx 1.173$ for the stiffer shell material $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$, see Fig. 3(d). The procedure demonstrates that effective concealment of the hole works also on this atomistic scale. As expected, the necessary thickness of the shell is larger for the softer shell material.
Quite remarkably, as shown in Fig. 4, results from atomistic simulations match well the results from analytical continuum elasticity theory over a wide range of ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for the considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34. This suggests that the picture provided by continuum elasticity theory remains valid even down to atomistic scales for the considered discrete systems of particles interacting solely via LJ pair potentials. As we gradually reduce the stiffness of the shell surrounding the hole in terms of the ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in Fig. 4, theory and simulations follow the same overall trend. Yet, quantitative differences appear in the necessary thickness of the shell around the hole that is necessary for effective concealment, as measured by $b/a$. One possible source for these deviations may be found in a progressively pronounced inhomogeneity on the atomic scale, in relation to the finite size of the simulation box. Indeed, the number of atoms within the shell, interacting via stronger interatomic forces compared to the background solid, grows from 0.87 $\%$ to 7.26 $\%$ of the total number of atoms in the solid with lower ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ [34]. Elevated thickness of the shell is necessary to conceal the hole with softer shell materials (lower $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$). Simultaneously, for fixed $a$ (here $20$ Å), elevated $b/a$ implies an increasing outer perimeter of the shell. This implies a larger number of particles interacting inhomogeneously, that is, with different LJ parameters concerning their interaction partners [34]. Conversely, for stiffer and thus thinner shells (higher $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$), the results from continuum theory and atomistic simulations are almost identical.
Figure 4: Comparison between the results from continuum elasticity theory and MD simulations. For given materials of $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34, we vary the ratio of the elastic moduli between the concealing shell around the enclosed hole and the surrounding elastic solid $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. We determine the ratio between the radius of the stiff shell $b$ that is necessary to mechanically conceal the presence of the enclosed hole of radius $a$ under uniform compression. As a result, we find the same qualitative trend of $b/a$ and quantitative agreement for stiff shells, that is, at elevated $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. (In the MD simulations, we set $a=$ 20 Å, implying a ratio between $a$ and box size $L$ of $a/L=0.089$.)
We remark that further quantitative agreement between analytical results and those from MD simulations can be achieved by reducing the ratio $a/L$ between the radius of the hole and the size of the simulated system. A corresponding demonstration for $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ is shown in Ref. 34. Yet, generally, if we reduce the size $a$ of the hole, the deviation in bulk elastic modulus is not that large already from the start. In Fig. 3(a) and (b) the red and black curves would be very close to each other, and so would be the blue curve for concealment. Finally, concerning actual parameter values of real materials, a hollow Al shell surrounded by a background solid of epoxy resin would have $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ within the range of 0.3–0.35, and $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}\approx 10$.
Conclusions. Summarizing, we considered a cylindrical hole that is surrounded by a stiffer cylindrical shell in a block of elastic material. The material is loaded uniformly under plane-strain conditions. Our scope was to identify for given material parameters a thickness of the shell that mechanically conceals the presence of the hole on the macroscopic level. In other words, the material as a whole, including such masked holes, behaves mechanically in the same way as if the holes were absent. Such a situation allows reduction in weight of components during light-weight construction, maintaining their overall shape and mechanical properties.
Indeed, a corresponding expression for the thickness of the shell around the hole was identified and derived as a function of the mechanical material parameters and the radius of the concealed hole. For this purpose, we assumed continuous, linearly elastic, homogeneous, isotropic materials under plane-strain conditions. The concentrations of the holes were low enough so that mechanical interactions between the holes can be neglected. One step further, we extended this possibility of mechanical concealment from a macroscopic continuum perspective down to discrete atomistic scales using molecular dynamics simulations. A purely atomistic, planar solid was considered. Consistent results were found. In principle, they validate the predictions of continuum elasticity theory down to the atomistic level. Gradual deviations in our finite-sized atomistic systems emerge at lower shell stiffnesses, which necessitate thicker shells. They imply a larger fraction of the stiffer shell component of the overall system, reducing the relative fraction of the surrounding elastic solid, and pronouncing the role of inhomogeneities.
Future extensions of this work are manifold, important, and obvious. An upcoming investigation shall address more general types of loading, beyond uniform. Besides plane-strain, also plane-stress and three-dimensional situations shall be considered in future works. Elevated concentrations of holes that make them mutually interact through their induced deformations shall be evaluated. More complex considerations include mechanically anisotropic materials such as crystalline solids. Altogether, we wish to support the further development of quantitative measures in the context of light-weight materials design.
Acknowledgments. We thank the European Union (EFRE) and the State Saxony Anhalt for support of this work through project no. ZS/2024/02/184030. Moreover, we thank Lukas Fischer, Johannes Menning, and Thomas Wallmersperger for stimulating discussions.